OpenStudy (anonymous):

Which of the following can be used to simplify the calculation of the expression 3 times forty-eight? 4 times 3 plus 8 times 3 the quantity of 3 times forty, end of quantity, plus 8 3 times the quantity of forty plus 8 3 times forty times 8

OpenStudy (domebotnos):

3(40 + 8) = 3*48 3(48) = 3*48 144 = 144 The answer is C.

OpenStudy (anonymous):

Nadine paid fourteen dollars and thirty cents for a bottle of lotion that was on sale at a "Buy One Get One Free" sale, so she got another one. How much did she pay, on average, per bottle? 4 dollars and seventy-seven cents 7 dollars and fifteen cents eleven dollars and ninety-two cents fourteen dollars and thirty cents

OpenStudy (domebotnos):

Post it in another question pls.